# In search of maximal branes on hyper-Kähler manifolds

### v1

@@ 3.2. Integral basis
Thanks to the work of Li–Qin–Wang [35, Theorem 1.1] and [51, Theorem 1.1], for a projective surface with vanishing odd Betti numbers, we have a concrete integral basis for the cohomology of its punctual Hilbert schemes.
@@ 3.3. Integral generators
We will need the following set of integral generators for the cohomology of Hilbert schemes by Li and Qin [35, Theorem 1.2].
@@ 4.1. Basics on compact hyper-Kähler manifolds
Given two compact hyper-Kähler manifolds $X,X^{\prime}$ that are deformation equivalent, a ring isomorphism $\phi\colon H^{*}(X,\mathbb{Z})\to H^{*}(X^{\prime},\mathbb{Z})$ is called a parallel transport operator, if there exist a smooth and proper family $\pi\colon\mathscr{X}\to B$ of compact hyper-Kähler manifolds over an analytic base $B$, two points $b,b^{\prime}\in B$ with isomorphisms $\psi\colon X\xrightarrow{\simeq}\mathscr{X}_{b}$ and $\psi^{\prime}\colon X^{\prime}\xrightarrow{\simeq}\mathscr{X}_{b^{\prime}}$, and a continuous path $\gamma$ from $b$ to $b^{\prime}$ such that $(\psi^{\prime})^{*}\circ\gamma_{*}\circ(\psi^{*})^{-1}=\phi$, where $\gamma_{*}$ is the parallel transport in the local system $\mathsf{R}\pi_{*}\mathbb{Z}$; see [41, Definition 1.1].
@@ 4.3. The monodromy group of manifolds of K3[n]-type
For a K3 surface $S$, we have the following computation of its monodromy group; see for example [27, Chapter 7, Proposition 5.5]:
@@
Let $X$ be a K3[n]-type hyper-Kähler manifold. Thanks to [42, Lemma 2.1], the natural morphism $\operatorname{Mon}(X)\to\operatorname{Mon}^{2}(X)$ is an isomorphism. From results of Markman [40, Theorem 1.2 and Lemma 4.2] we know that
@@
Markman proved in [39, Lemma 4.11] (combined with Lemma 4.10 in loc. cit.) that the Zariski closure of the subgroup $\operatorname{Mon}(X)\subset\operatorname{GL}(H^{*}(X,\mathbb{C}))$ is isomorphic to $O(H^{2}(X,\mathbb{C}))\times\{\pm 1\}$ if $X$ is of K3[n]-type with $n\geq 3$, and isomorphic to $O(H^{2}(X,\mathbb{C}))$ if $X$ is of K3[2]-type [^4] The K3[2]-type case is not stated in [39, Lemma 4.11], but it follows from the fact that in this case, the discriminant group $A_{H^{2}(X,\mathbb{Z})}$ of $H^{2}(X,\mathbb{Z})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ equipped with a non-zero quadratic form, hence $O(A_{H^{2}(X,\mathbb{Z})})$ is trivial. In particular, the discriminant character $\tau$ is trivial.. By this result of Zariski closure, we obtain that for any $n\geq 1$ and $X$ of K3[n]-type, a linear representation
@@ 4.4. Monodromy representation on the degree-4 cohomology
We need the following result of Markman [40, Theorem 1.10] on $Q(X,\mathbb{Z})$. Let $\overline{c}_{2}(X)$ denote the image of the second Chern class $c_{2}(X)$ under the natural projection $H^{4}(X,\mathbb{Z})\to Q(X,\mathbb{Z})$.
@@
In the statement of [40, Theorem 1.10], it was only indicated that there is some non-trivial character $\tau\colon\operatorname{Mon}^{2}(X)\to\{\pm 1\}$ making the diagram (4.6) commutative, without determining $\tau$ explicitly. However, it is easy to see that $\tau$ must be the discriminant character: since $Q(X,\mathbb{Z})$ is a unimodular lattice and $\sigma_{Q}(\overline{c}_{2}(X))=\overline{c}_{2}(X)$, $\sigma_{Q}$ must act trivially on the discriminant of the sublattice $\overline{c}_{2}(X)^{\perp}$ in $Q(X,\mathbb{Z})$. Therefore the transported action of $\sigma_{Q}$ to $H^{2}(X,\mathbb{Z})$ by conjugating with the isomorphism $e$ must act trivially on the discriminant of $H^{2}(X,\mathbb{Z})$. This implies that $\tau$ must be the discriminant character.
@@ 4.5. Monodromy representation for Hilbert schemes of points on K3 surfaces
We relate the action of monodromy operators of a K3 surface with the integral basis of the cohomology of the Hilbert schemes described in the previous sections. The results are based on Markman’s results and are collected in Oberdieck’s work [49, Section 3.6]. We briefly recall them for the reader’s convenience.
@@
Corollary 4.7 (Property 3 of [49, Section 3.6]).
@@
The following result will be useful to understand the action of the action of a monodromy operator that acts as $\operatorname{id}$ on the discriminant group $A_{H^{2}(X,\mathbb{Z})}$.
Lemma 4.9 (Property 2 of [49, Section 3.6]).
@@
$\rho_{n}(\operatorname{id}_{H^{2}(S^{[n]},\mathbb{C})},-1)=D\circ\rho_{n}(-\operatorname{id}_{H^{2}(S^{[n]},\mathbb{C})1},1),$
@@ 5.1. Topological constraints on free involutions of surfaces
Assume that $M$ has a CW-complex structure such that $\sigma$ is cellular. Consider the associated Smith–Gysin long exact sequence in the following form (see [11, Theorem 1.2.1] for example):
@@
The case where $n=1$ is exactly Theorem˜5.4. Assume $n\geq 2$ in the sequel. The assumption $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$ implies that $H^{1}(S,\mathbb{Z})=0$ and $H^{2}(S,\mathbb{Z})$ is 2-torsion-free. We have the following isomorphism (see for example [2, P.768]), where $(-)_{\operatorname{tf}}$ stands for $(-)/\operatorname{tors}$.
@@
Since $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, the cohomology $H^{*}(S,\mathbb{Z})$ is 2-torsion-free. By [55, Theorem 3.1 and the remark that follows], $H^{*}(S^{[n]},\mathbb{Z})$ is also 2-torsion-free. We obtain from (5.13) a $(\sigma^{[n]},\sigma)$-equivariant isomorphism:
@@
$\quad\quad\widetilde{\beta}$ $:=\mathfrak{p}_{-1}(\operatorname{pt})^{n-1}(\alpha)=(r^{*}(\operatorname{pt}))_{*}(\beta)\in H^{4n-2}(S^{[n]},\operatorname{\mathbb{F}_{2}}2).$
@@ 5.2. Kalinin spectral sequence
    has the following chain description (see [11, P.9]): for $x_{r}\in H_{r}(M,\operatorname{\mathbb{F}_{2}}2)^{\sigma}/\mathrm{Im}(1+\sigma_{*})$, we choose an $r$-dimensional cycle $\eta_{r}$ representing $x_{r}$. By assumption we can find an $(r+1)$-dimensional chain $\eta_{r+1}$ and an $(r+2)$-dimensional chain $\eta_{r+2}$, such that $\partial\eta_{r+2}=\eta_{r+1}+\sigma_{*}\eta_{r+1}$ and $\partial\eta_{r+1}=\eta_{r}+\sigma_{*}\eta_{r}$, then we define $d_{3}(x_{r})$ as the class of $\eta_{r+2}+\sigma_{*}\eta_{r+2}$. One can check that $d_{3}$ is well-defined in this case.
@@ 5.3. Quadrics and Hirzebruch surfaces
We use the notation in ˜5.9. Let e<0-e<0 be the self-intersection number of the exceptional section, which must be an even number since $\sigma$ has no fixed points. By [8] (see also [53] and [12, 2.5.2]), the real structure is actually unique up to isomorphism. By Degtyarev–Kharlamov [12], real rational surfaces are quasi-simple, hence the $\mathbb{R}$-surface $(S,\sigma)$ is real deformation equivalent to the $\mathbb{R}$-surface $(\mathbb{P}^{1}\times\mathbb{P}^{1},\tau\times\tau)$ where $\tau$ is the antipode real strucutre on $\mathbb{P}^{1}$; see [12, Section 4.2, Case 3] for the explicit construction of the deformation. Therefore, the Hilbert scheme $(S^{[n]},\sigma^{[n]})$ is real deformation equivalent to $((\mathbb{P}^{1}\times\mathbb{P}^{1})^{[n]},(\tau\times\tau)^{[n]})$. Since the latter is not maximal by Theorem˜5.11, while maximality being clearly a real deformation invariant property, we can conclude the non-maximality of $\sigma^{[n]}$. ∎
@@ 6.1. Natural anti-holomorphic involutions on Hilbert schemes
The following theorem completes Fu [17, Theorem 8.1] to an if-and-only-if result and generalizes Kharlamov–Răsdeaconu [32, Theorem 1.1, Theorem 1.2] to all dimensions. Even for $n=2$, our proof is different ([32] studies in much more details the fixed locus in the $n=2$ case).
@@
First, by [17, Proposition 4.4], for a maximal smooth projective surface $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, the connectedness of $S(\mathbb{R})$ is equivalent to $H^{2}(S,\mathbb{Z})^{\sigma}=0$, which in its turn is equivalent to the condition that $\sigma$ acts as $\operatorname{id}$ on $H^{2}(S,\mathbb{Z})$. The "if" part of the theorem is proven in Fu [17, Theorem 8.1]. For the "only if" part, let $\sigma$ denote both the associated anti-holomorphic involutions on $S$ and on $S^{[n]}$. Let $G$ be the cyclic group of order 2 generated by $\sigma$.
@@
For the connectedness of the real locus, let us assume that $S(\mathbb{R})$ is not connected, and show that $S^{[n]}$ is not maximal. By [17, Proposition 4.4], the non-connectedness of $S(\mathbb{R})$ implies that $H^{2}(S,\mathbb{Z})^{\sigma}\neq 0$. Choose a primitive element $\alpha\in H^{2}(S,\mathbb{Z})^{\sigma}$, and put
@@ 6.2. Natural holomorphic involutions on Hilbert schemes
where the last equality is by our assumption. Since $H^{2}_{G}(-,\mathbb{Z})$ classifies $G$-equivariant $\mathcal{C}^{\infty}$ complex line bundles (via the equivariant first Chern class map), we conclude that for any $\alpha\in H^{2}(S,\mathbb{Z})$, there exists a $G$-equivariant $\mathcal{C}^{\infty}$ complex line bundle $L_{\alpha}$ on $S$ with $c_{1}^{G}(L_{\alpha})=\alpha$. By [35, Proof of Lemma 3.5], we have :
@@ 6.3. Examples
Let us turn to some constructions of non-maximal involutions. We first point out the following application to hyper-Kähler geometry. It will be generalized in Section˜7 (see Theorem˜1.2, ˜1.3 and Theorem˜1.4).
@@
We first reduce the problem to the projective case. If $\sigma$ is holomorphic anti-symplectic, then $S$ is projective (see [46, Theorem 0.1]). If $\sigma$ is holomorphic symplectic or anti-holomorphic, then $S$ might not be projective but it is known that it can be deformed together with the involution to a projective K3 surface. For holomorphic symplectic involutions, it follows from the connectedness of their moduli space (obtained by Nikulin in loc. cit. and existence of such involutions on, say, the Fermat quartic in $\mathbb{P}^{3}$). For anti-holomorphic involutions, their moduli space is disconnected, but a similar approach can be applied. Namely, by [11, Theorem 13.8.1], the deformation type of a real K3 surface depends only on the isomorphism class of the lattice $H^{2}(S,\mathbb{Z})$ with an involution. Thus, it is sufficient to notice that the list of these isomorphism classes (summarized, for example, in [11, Theorem 8.4.2]) coincides with Nikulin’s list [47] for those classes that are realized by real quartic K3 surfaces in $\mathbb{P}^{3}$ (or to notice that, as it follows from [52], the number of classes realized by real K3 surfaces obtained as a double plane branched in a real curve of degree 6 is not smaller). Thus, as maximality is a deformation-invariant notion, we can assume $S$ to be projective in all the cases, and can apply Theorem˜6.3 and Theorem˜6.2.
@@
If $\sigma$ is anti-holomorphic, then by Theorem˜6.2, the assumption that $S^{[n]}$ is maximal implies that $S$ is maximal with connected real locus, or equivalently, by [17, Proposition 4.4], the whole $H^{2}(S,\mathbb{Z})$ is $\sigma$-anti-invariant. Now we have two ways to conclude: either use the fact that any maximal K3 surface has disconnected real locus by [11, Theorem 8.4.1] (see [32, Lemma 5.5] for a generalization), or argue that by ˜4.1, up to changing the complex structure on $S$ by a hyper-Kähler rotation, $\sigma$ becomes a holomorphic anti-symplectic involution, hence it must preserve the Kähler cone and cannot be $\operatorname{id}$ on $H^{2}(S,\mathbb{R})$. ∎
In fact, the statement for anti-holomorphic involutions in ˜6.6 holds more generally for surfaces with non-vanishing geometric genus. See the following generalization of [32, Corollary 1.3].
@@
Thanks to ˜6.8, starting from a smooth projective surface $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$ together with a maximal holomorphic or anti-holomorphic involution $\sigma$, after a successive blow-ups along fixed points of (lifted) involutions, we get a new surface $S^{\prime}$ equipped with involution $\sigma^{\prime}$, then the induced involution on the $n$-th Hilbert schemes of points on $S^{\prime}$ is maximal for $n\geq 2$ if and only if the same holds for the Hilbert schemes of $S$. In this way, from ˜6.4, ˜6.6 and ˜6.7, we can produce many maximal and non-maximal involutions on Hilbert schemes of surfaces. This generalizes [32, Corollary 1.4] to arbitrarily higher dimensions.
7. Nonexistence of maximal branes on $\operatorname{\mathrm{K3}}3^{[n]}$-type hyper-Kähler manifolds
In this section, we focus on the existence problem of maximal (holomorphic as well as anti-holomorphic) involutions on compact hyper-Kähler manifolds. In contrast to ˜6.6, we deal more generally with hyper-Kähler manifolds that are only deformation equivalent to Hilbert schemes of K3 surfaces, and involutions beyond the naturally induced ones. The main results have been stated in Introduction as Theorem˜1.2, ˜1.3 and Theorem˜1.4.

### v2

@@ 3.2. Integral basis
Thanks to the work of Li–Qin–Wang [41, Theorem 1.1] and [57, Theorem 1.1], for a projective surface with vanishing odd Betti numbers, we have a concrete integral basis for the cohomology of its punctual Hilbert schemes.
@@ 3.3. Integral generators
We will need the following set of integral generators for the cohomology of Hilbert schemes by Li and Qin [41, Theorem 1.2].
@@ 4.1. Basics on compact hyper-Kähler manifolds
Given two compact hyper-Kähler manifolds $X,X^{\prime}$ that are deformation equivalent, a ring isomorphism
$\phi\colon H^{*}(X,\mathbb{Z})\to H^{*}(X^{\prime},\mathbb{Z})$
is called a parallel transport operator, if there exist a smooth and proper family $\pi\colon\mathscr{X}\to B$ of compact hyper-Kähler manifolds over an analytic base $B$, two points $b,b^{\prime}\in B$ with isomorphisms $\psi\colon X\xrightarrow{\simeq}\mathscr{X}_{b}$ and $\psi^{\prime}\colon X^{\prime}\xrightarrow{\simeq}\mathscr{X}_{b^{\prime}}$, and a continuous path $\gamma$ from $b$ to $b^{\prime}$ such that $(\psi^{\prime})^{*}\circ\gamma_{*}\circ(\psi^{*})^{-1}=\phi$, where $\gamma_{*}$ is the parallel transport in the local system $\mathsf{R}\pi_{*}\mathbb{Z}$; see [47, Definition 1.1].
@@ 4.3. The monodromy group of manifolds of K3[n]-type
For a K3 surface $S$, we have the following computation of its monodromy group; see for example [31, Chapter 7, Proposition 5.5]:
@@
Let $X$ be a K3[n]-type hyper-Kähler manifold. Thanks to [48, Lemma 2.1], the natural morphism $\operatorname{Mon}(X)\to\operatorname{Mon}^{2}(X)$ is an isomorphism. From results of Markman [46, Theorem 1.2 and Lemma 4.2] we know that
@@
Markman proved in [45, Lemma 4.11] (combined with Lemma 4.10 in loc. cit.) that the Zariski closure of the subgroup $\operatorname{Mon}(X)\subset\operatorname{GL}(H^{*}(X,\mathbb{C}))$ is isomorphic to $O(H^{2}(X,\mathbb{C}))\times\{\pm 1\}$ if $X$ is of K3[n]-type with $n\geq 3$, and isomorphic to $O(H^{2}(X,\mathbb{C}))$ if $X$ is of K3[2]-type [^4] The K3[2]-type case is not stated in [45, Lemma 4.11], but it follows from the fact that in this case, the discriminant group $A_{H^{2}(X,\mathbb{Z})}$ of $H^{2}(X,\mathbb{Z})$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$ equipped with a non-zero quadratic form, hence $O(A_{H^{2}(X,\mathbb{Z})})$ is trivial. In particular, the discriminant character $\tau$ is trivial.. By this result of Zariski closure, we obtain that for any $n\geq 1$ and $X$ of K3[n]-type, a linear representation
@@ 4.4. Monodromy representation on the degree-4 cohomology
We need the following result of Markman [46, Theorem 1.10] on $Q(X,\mathbb{Z})$. Let $\overline{c}_{2}(X)$ denote the image of the second Chern class $c_{2}(X)$ under the natural projection $H^{4}(X,\mathbb{Z})\to Q(X,\mathbb{Z})$.
@@
In the statement of [46, Theorem 1.10], it was only indicated that there is some non-trivial character $\tau\colon\operatorname{Mon}^{2}(X)\to\{\pm 1\}$ making the diagram (4.6) commutative, without determining $\tau$ explicitly. However, it is easy to see that $\tau$ must be the discriminant character: since $Q(X,\mathbb{Z})$ is a unimodular lattice and $\sigma_{Q}(\overline{c}_{2}(X))=\overline{c}_{2}(X)$, $\sigma_{Q}$ must act trivially on the discriminant of the sublattice $\overline{c}_{2}(X)^{\perp}$ in $Q(X,\mathbb{Z})$. Therefore the transported action of $\sigma_{Q}$ to $H^{2}(X,\mathbb{Z})$ by conjugating with the isomorphism $e$ must act trivially on the discriminant of $H^{2}(X,\mathbb{Z})$. This implies that $\tau$ must be the discriminant character.
@@ 4.5. Monodromy representation for Hilbert schemes of points on K3 surfaces
We relate the action of monodromy operators of a K3 surface with the integral basis of the cohomology of the Hilbert schemes described in the previous sections. The results are based on Markman’s results and are collected in Oberdieck’s work [55, Section 3.6]. We briefly recall them for the reader’s convenience.
@@
Corollary 4.7 (Property 3 of [55, Section 3.6]).
@@
The following result will be useful to understand the action of a monodromy operator that acts as $\operatorname{id}$ on the discriminant group $A_{H^{2}(X,\mathbb{Z})}$.
Lemma 4.9 (Property 2 of [55, Section 3.6]).
@@
$\rho_{n}(\operatorname{id}_{H^{2}(S^{[n]},\mathbb{C})},-1)=D\circ\rho_{n}(-\operatorname{id}_{H^{2}(S^{[n]},\mathbb{C})},1),$
@@ 5.1. Topological constraints on free involutions of surfaces
Assume that $M$ has a CW-complex structure such that $\sigma$ is cellular. Consider the associated Smith–Gysin long exact sequence in the following form (see [12, Theorem 1.2.1] for example):
@@
The case where $n=1$ is exactly Theorem˜5.4. Assume $n\geq 2$ in the sequel. The assumption $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$ implies that $H^{1}(S,\mathbb{Z})=0$ and $H^{2}(S,\mathbb{Z})$ is 2-torsion-free. We have the following isomorphism (see for example [3, P.768]), where $(-)_{\operatorname{tf}}$ stands for $(-)/\operatorname{tors}$.
@@
Since $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, the cohomology $H^{*}(S,\mathbb{Z})$ is 2-torsion-free. By [61, Theorem 3.1 and the remark that follows], $H^{*}(S^{[n]},\mathbb{Z})$ is also 2-torsion-free. We obtain from (5.13) a $(\sigma^{[n]},\sigma)$-equivariant isomorphism:
@@
$\qquad\qquad\widetilde{\beta}$ $:=\mathfrak{p}_{-1}(\operatorname{pt})^{n-1}(\alpha)=(r^{*}(\operatorname{pt}))_{*}(\beta)\in H^{4n-2}(S^{[n]},\operatorname{\mathbb{F}_{2}}2).$
@@ 5.2. Kalinin spectral sequence
    has the following chain description (see [12, P.9]): for $x_{r}\in H_{r}(M,\operatorname{\mathbb{F}_{2}}2)^{\sigma}/\mathrm{Im}(1+\sigma_{*})$, we choose an $r$-dimensional cycle $\eta_{r}$ representing $x_{r}$. By assumption we can find an $(r+1)$-dimensional chain $\eta_{r+1}$ and an $(r+2)$-dimensional chain $\eta_{r+2}$, such that $\partial\eta_{r+2}=\eta_{r+1}+\sigma_{*}\eta_{r+1}$ and $\partial\eta_{r+1}=\eta_{r}+\sigma_{*}\eta_{r}$, then we define $d_{3}(x_{r})$ as the class of $\eta_{r+2}+\sigma_{*}\eta_{r+2}$. One can check that $d_{3}$ is well-defined in this case.
@@ 5.3. Quadrics and Hirzebruch surfaces
We use the notation in ˜5.9. Let e<0-e<0 be the self-intersection number of the exceptional section, which must be an even number since $\sigma$ has no fixed points. By [9] (see also [59] and [13, 2.5.2]), the real structure is actually unique up to isomorphism. By Degtyarev–Kharlamov [13], real rational surfaces are quasi-simple, hence the $\mathbb{R}$-surface $(S,\sigma)$ is real deformation equivalent to the $\mathbb{R}$-surface $(\mathbb{P}^{1}\times\mathbb{P}^{1},\tau\times\tau)$ where $\tau$ is the antipode real strucutre on $\mathbb{P}^{1}$; see [13, Section 4.2, Case 3] for the explicit construction of the deformation. Therefore, the Hilbert scheme $(S^{[n]},\sigma^{[n]})$ is real deformation equivalent to $((\mathbb{P}^{1}\times\mathbb{P}^{1})^{[n]},(\tau\times\tau)^{[n]})$. Since the latter is not maximal by Theorem˜5.11, while maximality being clearly a real deformation invariant property, we can conclude the non-maximality of $\sigma^{[n]}$. ∎
@@ 6.1. Natural anti-holomorphic involutions on Hilbert schemes
The following theorem completes Fu [20, Theorem 8.1] to an if-and-only-if result and generalizes Kharlamov–Răsdeaconu [39, Theorem 1.1, Theorem 1.2] to all dimensions. Even for $n=2$, our proof is different ([39] studies in much more details the fixed locus in the $n=2$ case).
@@
First, by [20, Proposition 4.4], for a maximal smooth projective surface $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$, the connectedness of $S(\mathbb{R})$ is equivalent to $H^{2}(S,\mathbb{Z})^{\sigma}=0$, which in its turn is equivalent to the condition that $\sigma$ acts as $\operatorname{id}$ on $H^{2}(S,\mathbb{Z})$. The "if" part of the theorem is proven in Fu [20, Theorem 8.1]. For the "only if" part, let $\sigma$ denote both the associated anti-holomorphic involutions on $S$ and on $S^{[n]}$. Let $G$ be the cyclic group of order 2 generated by $\sigma$.
@@
For the connectedness of the real locus, let us assume that $S(\mathbb{R})$ is not connected, and show that $S^{[n]}$ is not maximal. By [20, Proposition 4.4], the non-connectedness of $S(\mathbb{R})$ implies that $H^{2}(S,\mathbb{Z})^{\sigma}\neq 0$. Choose a primitive element $\alpha\in H^{2}(S,\mathbb{Z})^{\sigma}$, and put
@@ 6.2. Natural holomorphic involutions on Hilbert schemes
where the last equality is by our assumption. Since $H^{2}_{G}(-,\mathbb{Z})$ classifies $G$-equivariant $\mathcal{C}^{\infty}$ complex line bundles (via the equivariant first Chern class map), we conclude that for any $\alpha\in H^{2}(S,\mathbb{Z})$, there exists a $G$-equivariant $\mathcal{C}^{\infty}$ complex line bundle $L_{\alpha}$ on $S$ with $c_{1}^{G}(L_{\alpha})=\alpha$. By [41, Proof of Lemma 3.5], we have :
@@ 6.3. Examples
Let us turn to some constructions of non-maximal involutions. We first point out the following application to hyper-Kähler geometry. It will be generalized in Section˜7 (see Theorem˜1.4, ˜1.5 and Theorem˜1.6).
@@
We first reduce the problem to the projective case. If $\sigma$ is holomorphic anti-symplectic, then $S$ is projective (see [53, Theorem 0.1]). If $\sigma$ is holomorphic symplectic or anti-holomorphic, then $S$ might not be projective but it is known that it can be deformed together with the involution to a projective K3 surface. For holomorphic symplectic involutions, it follows from the connectedness of their moduli space (obtained by Nikulin in loc. cit. and existence of such involutions on, say, the Fermat quartic in $\mathbb{P}^{3}$). For anti-holomorphic involutions, their moduli space is disconnected, but a similar approach can be applied. Namely, by [12, Theorem 13.8.1], the deformation type of a real K3 surface depends only on the isomorphism class of the lattice $H^{2}(S,\mathbb{Z})$ with an involution. Thus, it is sufficient to notice that the list of these isomorphism classes (summarized, for example, in [12, Theorem 8.4.2]) coincides with Nikulin’s list [54] for those classes that are realized by real quartic K3 surfaces in $\mathbb{P}^{3}$ (or to notice that, as it follows from [58], the number of classes realized by real K3 surfaces obtained as a double plane branched in a real curve of degree 6 is not smaller). Thus, as maximality is a deformation-invariant notion, we can assume $S$ to be projective in all the cases, and can apply Theorem˜6.3 and Theorem˜6.2.
@@
If $\sigma$ is anti-holomorphic, then by Theorem˜6.2, the assumption that $S^{[n]}$ is maximal implies that $S$ is maximal with connected real locus, or equivalently, by [20, Proposition 4.4], the whole $H^{2}(S,\mathbb{Z})$ is $\sigma$-anti-invariant. Now we have two ways to conclude: either use the fact that any maximal K3 surface has disconnected real locus by [12, Theorem 8.4.1] (see [39, Lemma 5.5] for a generalization), or argue that by ˜4.1, up to changing the complex structure on $S$ by a hyper-Kähler rotation, $\sigma$ becomes a holomorphic anti-symplectic involution, hence it must preserve the Kähler cone and cannot be $\operatorname{id}$ on $H^{2}(S,\mathbb{R})$. ∎
In fact, the statement for anti-holomorphic involutions in ˜6.6 holds more generally for surfaces with non-vanishing geometric genus. See the following generalization of [39, Corollary 1.3].
@@
Thanks to ˜6.8, starting from a smooth projective surface $S$ with $H^{1}(S,\operatorname{\mathbb{F}_{2}}2)=0$ together with a maximal holomorphic or anti-holomorphic involution $\sigma$, after a successive blow-ups along fixed points of (lifted) involutions, we get a new surface $S^{\prime}$ equipped with involution $\sigma^{\prime}$, then the induced involution on the $n$-th Hilbert schemes of points on $S^{\prime}$ is maximal for $n\geq 2$ if and only if the same holds for the Hilbert schemes of $S$. In this way, from ˜6.4, ˜6.6 and ˜6.7, we can produce many maximal and non-maximal involutions on Hilbert schemes of surfaces. This generalizes [39, Corollary 1.4] to arbitrarily higher dimensions.
7. Nonexistence of maximal brane involutions on hyper-Kähler manifolds of $\operatorname{\mathrm{K3}}3^{[n]}$-type
In this section, we focus on the existence problem of maximal (holomorphic as well as anti-holomorphic) involutions on compact hyper-Kähler manifolds. In contrast to ˜6.6, we deal more generally with hyper-Kähler manifolds that are only deformation equivalent to Hilbert schemes of K3 surfaces, and involutions beyond the naturally induced ones. The main results have been stated in Introduction as Theorem˜1.4, ˜1.5 and Theorem˜1.6.